OpenStudy (christian_10_):

Find the solution for the equation 5(x-4)-4x= 18-23 Fan & medal. :)

OpenStudy (ttp):

Hold up!!

OpenStudy (christian_10_):

Ok xD

OpenStudy (samigupta8):

5x-20-4x=-5 Now....x-20=-5 Then x=20-5=15

imqwerty (imqwerty):

1 Simplify 18−23 to −5 5(x−4)−4x=−5 2 Expand 5x−20−4x=−5 3 Simplify 5x−20−4x to x−20 x−20=−5 4 Add 20 to both sides x=−5+20 5 Simplify −5+20 to 15 x=15